Translate the following into a quadratic equation, and solve it:
The length of a rectangular garden is two more than its width; if the area of the garden is 35 square meters, what are its dimensions?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Translate the following into a quadratic equation, and solve it:
The length of a rectangular garden is two more than its width; if the area of the garden is 35 square meters, what are its dimensions?
------
Let width be "x"; then length is "x+2".
Area = (length)(width)
35 = (x+2)(x)
x^2+2x-35 = 0
Factor:
(x+7)(x-5) = 0
Positive solution:
x = 5 meters (width)
x+2 = 7 meters (length)
===========================
Cheers,
Stan H.
---------------------

Answer by mananth(16949)

width =x
length = (x+2)
Area = x(x+2)= 35
x^2+2x-35=0
x^2+7x-5x-35=0
x(x+7)-5(x+7)=0
(x+7)(x-5)=0
width = 5m
length = 5+2=7m